# TV-TREES: Multimodal Entailment Trees for Neuro-Symbolic Video Reasoning

### v3

@@ 4.3 Evidence Retrieval
Existing natural language inference (NLI) models are not well-suited for classifying entailments within highly contextual and social dialogue, which often insinuates meaning not directly stated within the text. Instead of producing an entirely new dataset for the domain of dialogue NLI, we use GPT-3.5 to generate a set of natural language inferences about the dialogue sample written in the style as data points in a dataset akin to SNLI Bowman et al. (2015b), conditioned on a question form of the hypothesis, $q$. Presenting the question under discussion in the interrogative form significantly reduces the hallucination rate compared to passing in the original hypothesis. $q$ is also generated via GPT-3.5 taking the hypothesis $h$ as input.
Our system queries GPT for five inferences from a given question and passage. Then, we run these inferences through GPT to verify that they are entailed by the transcript. Examples of generated inferences are included in Figure 3.
4.4 Evidence Filtering
@@
We use a cross-encoder trained on SNLI and MultiNLI to determine whether any of the retrieved evidence sufficiently entails the hypothesis. We accept any sample that achieves a logit score above a certain threshold for the "entailment" label.
@@
Then, we apply a secondary entailment filter that ensures the inferences are accurate descriptions of the content presented in the dialogue. This is important as, while conditioning the inference generator on an interrogative form of the hypothesis mitigates hallucinations, it does not eliminate them entirely. Identifying these cases is attempted through a GPT filter that takes in the inference and the dialogue, without any hypothesis conditioning.
@@
Finally, as the cross-encoder tends to ignore negation, which is often present in the generated inferences, we additionally pass the filtered inference-hypothesis pairs to a GPT-3.5 prompt that verifies the entailment.
The system only retains the inferences that pass through all three filters.
4.5 Decomposition
@@
In the case where no atomic evidence can be retrieved from the transcript or video that immediately entails the current hypothesis, the system attempts to break it down into two sub-hypotheses that are (1) complete sentences without ambiguous pronouns or decontextualized references and (2) compositionally equivalent to the original hypothesis, i.e., proving the two sub-hypotheses as true is approximately logically equivalent to proving the original hypothesis.
We prompt GPT-3.5 to break the current hypothesis into two compositionally equivalent pieces of information, conditioned on the dialogue sample extracted in section 4.2. We instruct GPT to only return a decomposition when it is syntactically possible, to avoid recursing on sentence fragments and hypothesis repeats that the model may erroneously output if a sound decomposition cannot be found.
  Method Zero-Shot Full Clips Transparent Dialogue Vision TVQA Acc.
@@
STAGE No Yes No Yes Yes 70.5
HERO No No No Yes Yes 74.2
FrozenBiLM No No No Yes Yes 82.0
LLaMA-VQA No No No Yes Yes 82.2
@@
$*{}^{*}$ Yes Yes No Yes Yes 26.3
SeVILA Yes Yes No No Yes 38.2
VideoChat2 Yes Yes No No Yes 40.6
${}^{\text{\textdaggerdbl}}$ Yes Yes Yes Yes No 44.9
TV-TREES Yes Yes Yes Yes Yes 49.4
@@
Table 1: Table comparing various vision-text understanding models across a set of criteria including performance on the TVQA benchmark. All zero-shot methods (Zero-Shot) take in full video clips (Full Clips), but unlike the fine-tuned approaches, none except FrozenBiLM operate over both vision and dialogue modalities. Notably, TV-TREES is the only interpretable approach. Experiment results suggest that TV-TREES and TV-TREES with text input only (${}^{\text{\textdaggerdbl}}$) outperform existing zero-shot methods on full clips. All numbers for competing approaches are as they are reported in their respective papers except for FrozenBiLM*, which we re-run on our validation subset with full clips as input. (On ground truth clip fragments, FrozenBiLM reports 59.7% accuracy). Results suggest that a more robust visual understanding module could further improve the performance of TV-TREES, seeing the baseline results achieved by models taking in vision input only.
4.6 Visual Reasoning
We pass in the questions generated in Section 5.3 alongside video frames from the localized evidence window (if applicable) sampled at 2 FPS into a vision-language model. In our experiments, we use LLaVA-7B Liu et al. (2023). To encourage conservative classifications, in addition to asking for “yes" and “no" answers we encourage the model to respond with “not enough information" if it is unsure or the image does not provide sufficient evidence. If more than 10% of the frames in the window result in an affirmative answer from the VLM model, the visual content is considered to contain sufficient entailing evidence, and the frame with the highest logits score is returned. If no frames result in an affirmative answer, no appropriate visual evidence entails the hypothesis. The LLaVA-7B prompt is included alongside the GPT prompts in Appendix A.
We also use GPT-3.5 to anonymize the question generated in section 4.1, replacing character names with common nouns such as "person". We query LLaVA-7B on each frame individually, using the anonymized question as textual input. We compare the performance of this approach to providing the original question in Appendix B, but find that the modification makes a marginal difference (approximately one point lower performance on average).
5 Evaluation Methodology
Traditionally, qualitative natural text evaluations have often been conducted using humans Celikyilmaz et al. (2021), either expert annotators or crowdsourced workers. Recently, researchers have considered whether these human evaluations could be replaced by high-performing LLMs like GPT-4 Naismith et al. (2023). Following this line of thinking, in this section, we detail how we implement the evaluation metrics described in Section 3.2 through human annotations as well as GPT-4. We report evaluation statistics for both methods in Section 6.
5.1 Human Evaluations
Considering the three evaluation metrics described in Section 3.2 (acceptability, relevance, and sufficiency), we evaluate trees along these qualia through three annotation tasks. The first task provides annotators with the visual or text evidence assigned to the leaf nodes by the algorithm and asks them to assess the correctness of the leaf node hypotheses on a scale of 1-5 (acceptability) based on that evidence. The second task provides annotators with $(h_{0},h^{\prime})$ pairs from branching nodes and asks if the child hypothesis $h^{\prime}$ is relevant to the parent $h_{0}$ (relevance). The third task provides annotators with a full hypothesis triplet $(h_{0},h_{1},h_{2})$ from a branching node with parent $h_{0}$ and child premises $h_{1}$ and $h_{2}$ and asks (1) whether $h_{1}$ and $h_{2}$ each introduce distinct information (the other facet of relevance, we also call this distinctness for disambiguation purposes), and (2) if $h_{0}$ introduces information not provided by $h_{1}$ and $h_{2}$ together, to check for entailment (sufficiency). Through these tasks, annotators are also asked to indicate if any of the hypotheses or premises are malformed or otherwise uninterpretable (the other facet of acceptability).
Every node in a multimodal entailment tree is assigned a binary score for each assessment described above (except for the correctness checks, which are collected on a scale of 1-5). We include all task instructions and layouts in Appendix D, along with more formal descriptions of the five quantitative acceptability scores.
5.2 GPT Evaluations
We take the qualia outlined in Section 3.2 and write three GPT-4 prompts for (1) correct leaves in the text domain, (2) correct leaves in the vision domain, and (3) the remaining three checklist items. Correctness check prompts are modality dependent as we use GPT-4V for vision evaluations, and separate both from the remaining checklist items as only the leaves must be evaluated for evidence-centric correctness. We use the same scoring values as in the human evaluations and pass in twelve decompositions per prompt for the text prompts. These prompts are included in full in Appendix E.
5.3 Tree Scoring Paradigm
We consider the mean normalized score of the three main evaluation qualia across all nodes as the overall “composition score" for each individual tree:
$S=\frac{a+s+0.5(d+r)}{3}$
where $a$ is the tree’s mean leaf acceptability score, $d$ is the tree’s mean distinctness score, $r$ is the tree’s mean relevance score, and $s$ is the tree’s mean sufficiency score.
  Trees Acceptability Relevance Distinctness Sufficiency Score
@@ 5.3 Tree Scoring Paradigm
Table 2: Entailment tree quality evaluations using human and LLM evaluators. For the human annotations, acceptability corresponds to Task 1, relevance to Task 2, and distinctness and sufficiency to Task 3. These metrics are explicitly labeled in the GPT-4 prompts for evaluation. This table reports mean scores aggregated per tree for each category. In addition to metric scores, we report composition score as defined in Section 5.3. We partition results by modality: We report scores for trees using text content only, trees that use visual evidence, and both groups combined. As shown, tree scores largely suffer due to the correctness of the leaf nodes, which is unsurprising given the difficulty of extracting high-level inferences from social dialogue and often ambiguous video screenshots.
  Method Acc. Comp. Acc. Comp. %
Vision 32.4 51.9 19.7
Dialogue 44.9 53.3 51.5
Both 49.4 53.0 69.5
Table 3: Ablation experiment results comparing performance on TVQA when using only dialogue evidence, only visual evidence, and both modalities as evidence. We report the overall accuracy, the accuracy of the system on questions where at least one proof was complete, and the percentage of questions on which at least one proof was complete.
6 Experiments
We evaluate TV-TREES on the TVQA dataset, comparing its performance against a text-only version of the architecture and competing zero-shot VideoQA approaches. We compare all approaches in terms of QA accuracy and compare the entailment tree generation methods in terms of tree quality as described in Section 5.
6.1 Setup
We instantiate TV-TREES as it is described in Section 4, setting the maximum recursion depth to $k=2$, or allowing trees with up to 3 levels. Our experiments focus on the multiple-choice VideoQA domain, and so we consider a question’s correct answer to be the answer that results in a complete tree. In the case that the system does not successfully complete any tree for the five answer candidates, we consider the answer candidate with the "most complete" tree to be the correct answer, breaking ties by the average entailment score at each node. When complete trees are generated for multiple answers, we break ties in the same way.
6.2 Evaluation on TVQA
Data   We evaluate our system on 3,000 multiple choice questions from the validation set of TVQA Lei et al. (2018). TVQA is a VideoQA benchmark that includes multiple choice questions about the dialogue and visual content of video clips taken from six TV shows. The clips are approximately 60-90 seconds long and contain around 30 lines of dialogue each. An example TVQA question is shown in Figure 1.
Models   In the zero-shot setting, in addition to TV-TREES, we consider zero-shot approaches FrozenBiLM Yang et al. (2022), SeVILA Yu et al. (2023), and VideoChat2 Li et al. (2023a). We also include performance reported by other systems (not zero-shot) for context: STAGE Lei et al. (2019), HERO Li et al. (2020), FrozenBiLM (fine-tuned) Yang et al. (2022), and LLaMA-VQA Ko et al. (2023).
Ablations   Existing work notes that both existing multimodal models are biased toward the text modality, often relying on text data for reasoning even for video-centric questions. In line with this theme, we evaluate our system’s performance conditioned on input modality on a subset of the TVQA validation set. We first evaluate the system when it is only provided with dialogue transcripts from the clip and then when it is only provided with video frames from the clip.
Results   We report overall accuracy alongside qualitative comparisons between the approaches in Table 4. As shown in the table, TV-TREES outperforms existing zero-shot methods when using full clips, but still shows significant room for future improvements. Notably, the text-only model outperforms joint-modality methods, and the joint-modality model only improves performance modestly, suggesting that the language modules of TV-TREES are more robust and performance could be further increased through improvements to the vision pipeline. This is further shown in the ablation experiment results in Table 3, which suggests that vision evidence alone allows TV-TREES to complete trees for only 19.7% of the questions compared to 51.5% and 69.5% for text-only and joint-modality models, respectively.
6.3 Proof Scoring
Setup   We randomly sample 600 completed entailment trees generated by TV-TREES on the TVQA validation split, split evenly between text-only and multimodal trees and split evenly among tree complexity (ranging from one to seven tree nodes).
We evaluate these sampled trees using the automatic GPT4 approach as described in Section 5.2. We then sample 200 proofs from this set (evenly distributed across modalities and complexity) and we annotate this set with human annotators from Amazon Mechanical Turk as described in Section 5.1. For human annotations, we identify careful annotators through a preliminary pilot task where each annotator’s work is scored by hand, and only high-scoring annotators are invited to annotate the full proofs. More information regarding these crowdsourced annotations is included in Appendix C.
For scoring acceptability, we provide the scorer with the localized dialogue retrieved by the cross-encoder model described in Section 4.2, and the video frames that achieved the highest logits scores during VQA inference, depending on the modality. We report results in Table 2.
Results   Generally, there is a close alignment between the GPT-4 and human scores. While the overall average score assigned to the trees is within a .1 point difference between the two approaches, GPT-4 tended to score the text-only trees more harshly than humans, and the multimodal trees more leniently. This is shown primarily in the resulting acceptability scores, and more moderately in the sufficiency scores. GPT-4 rated relevance scores more leniently for both modalities, which may stem from differences in human interpretations of the task instructions. In contrast, distinctness scores are almost identical between the two methods.
We find that, unsurprisingly, the majority of errors in the produced trees stem from acceptability issues. According to human evaluations, the visual module produces lower-quality inferences than the textual modules do. This is not surprising, as we are able to include additional entailment filters for the textual reasoning steps to remove lower-quality predictions before constructing the final entailment trees, whereas we do not have similar methods in place for visual inference. Based on these results, introducing stronger entailment classifiers for both domains may significantly improve performance on tree evaluation as well as on general VideoQA.

### v4

@@ 4.3 Evidence Retrieval (Retrieve)
Existing natural language inference (NLI) models are not well-suited for classifying entailments within highly contextual and social dialogue, which often requires sophisticated inferential ability. Therefore, we use GPT-3.5 to generate a set of five natural language inferences about the dialogue, conditioned on a question form of the hypothesis, $q$, written in the style of a dataset like SNLI Bowman et al. (2015b). Presenting the question under discussion in the interrogative form significantly reduces the hallucination rate compared to passing in the original hypothesis. $q$ is also generated via GPT-3.5 taking the hypothesis $h$ as input. Example generated inferences are shown in Figure 3.
4.4 Evidence Filtering (Filter)
@@
We use a cross-encoder trained on SNLI and MNLI Williams et al. (2017)333huggingface.co/cross-encoder/nli-distilroberta-base to determine whether any of the retrieved inferences entail the hypothesis. We accept any sample that achieves a logit score above a given threshold.
@@
We apply a secondary entailment filter using GPT-3.5 that ensures the inferences are accurate descriptions of the content presented in the dialogue. This is important because, while conditioning the inference generator on an interrogative form of the hypothesis mitigates hallucinations, it does not eliminate them entirely.
@@
Finally, as the cross-encoder tends to ignore negation, we additionally pass the filtered inference-hypothesis pairs to a GPT-3.5 prompt to verify the entailment a final time. The system retains the inferences that pass all three filters.
4.5 Decomposition (Decompose)
@@
If no retrieved evidence entails the current hypothesis, TV-TREES breaks down the hypothesis into two sub-hypotheses that are (1) complete sentences without ambiguous pronouns or decontextualized references and (2) compositionally equivalent to the original hypothesis, i.e., proving the two sub-hypotheses as true is approximately logically equivalent to proving the original hypothesis. We prompt GPT-3.5 to break the current hypothesis into two compositionally equivalent pieces of information.
4.6 Visual Reasoning (Filter)
We pass in the questions generated in Section 4.5 alongside video frames from the localized evidence window (if applicable) into a vision-language model. In our experiments, we use LLaVA-7B Liu et al. (2023). [^4] The LLaVA-7B prompt is included in Appendix A. To encourage conservative classifications, in addition to asking for “yes" and “no" answers we encourage the model to respond with “not enough information" if it is unsure. If more than 10% of the frames in the window result in an affirmative answer from the VLM model, the visual content is considered to contain sufficient entailing evidence, and the frame with the highest logits score is returned.
5 Evaluation Methodology
Traditionally, qualitative natural text evaluations have been conducted using humans Celikyilmaz et al. (2021). Recently, researchers have considered whether these human evaluations could be replaced by high-performing LLMs like GPT-4 Naismith et al. (2023). We detail how we implement the evaluation metrics described in Section 3.2 with both human annotators and GPT-4. We report evaluation results for both methods in Section 6.
5.1 Human Evaluations
We evaluate trees using the metrics introduced in Section 3.2 (acceptability, relevance, and sufficiency) through three annotation tasks. The first task provides annotators with a tree’s leaf node evidence (images or text) and asks them to assess the correctness of the leaf node hypotheses on a scale of 1-5 (acceptability) based on that evidence. The second task provides annotators with parent-child hypothesis $(h_{0},h^{\prime})$ pairs from branching nodes and asks if the child hypothesis $h^{\prime}$ is relevant to the parent $h_{0}$ (relevance). The third task provides annotators with a full hypothesis triplet $(h_{0},h_{1},h_{2})$ from a branching node with parent $h_{0}$ and child hypotheses $h_{1}$ and $h_{2}$ and asks (1) whether $h_{1}$ and $h_{2}$ each introduce distinct information (the other facet of relevance, we also call this distinctness for disambiguation purposes), and (2) if $h_{0}$ introduces information not provided by $h_{1}$ and $h_{2}$ together, to check for entailment (sufficiency). Through these tasks, annotators are also asked to indicate if any of the hypotheses or premises are malformed or otherwise uninterpretable (also acceptability).
Every node in a multimodal entailment tree is assigned a binary score for each assessment described above (except for the correctness checks, which are collected on a scale of 1-5). We include all task instructions and layouts in Appendix D.
5.2 GPT Evaluations
We take the qualia outlined in Section 3.2 and write three GPT-4 prompts testing (1) acceptability of evidence in the text domain, (2) acceptability of evidence in the vision domain, and (3) relevancy and sufficiency. [^5] These prompts are included in full in Appendix E. We use the same scoring values used in the human evaluations.
5.3 Tree Scoring Paradigm
We consider the mean normalized score of the three main evaluation qualia across all nodes as the overall “composition score" for each individual tree,
$S=\frac{1}{3}(a+s+0.5(d+r))$
where $a$ is the tree’s mean normalized acceptability score, $d$ is the mean distinctness score, $r$ is the mean relevance score, and $s$ is the mean sufficiency score.
6 Experiments
We evaluate TV-TREES using the TVQA dataset as input data. We compare its zero-shot QA performance against competing video-language QA approaches to illustrate its practical usage, evaluate its overall tree quality through our evaluation method described in Section 5, and organize its reasoning error modes through a qualitative study.
Setup
We instantiate TV-TREES as it is described in Section 4, allowing for trees with up to 3 levels ($k=2$). Our experiments focus on the multiple-choice QA domain, and so we consider a question’s correct answer to be the answer that results in a complete tree. In the case that the system does not successfully complete any tree for the five answer candidates, we consider the answer candidate with the "most complete" tree to be the correct answer, breaking ties by average entailment score. When complete trees are generated for multiple answers, we break ties in the same way.
6.1 QA Evaluation
We focus on video-language QA to take the first step adapting a text-only method to other domains, but complex video-language benchmarks are sparse: TVQA and MovieQA Tapaswi et al. (2016) are the two commonly used video-language datasets, but past research suggests that about half of MovieQA questions can be answered without reasoning over the video content Jasani et al. (2019). Therefore, we focus our study on TVQA.
Data
We evaluate our system on 3,000 multiple choice questions from the validation set of TVQA Lei et al. (2018). TVQA is a video-language QA benchmark that includes multiple choice questions about the dialogue and visual content of video clips taken from TV shows. The clips are about 60-90 seconds long and contain around 30 lines of dialogue. A sample question is shown in Figure 1.
Models
In the zero-shot setting, we consider zero-shot systems FrozenBiLM Yang et al. (2022), SeVILA Yu et al. (2023), and VideoChat2 Li et al. (2023a). We also include performance reported by other systems (not zero-shot) for context: STAGE Lei et al. (2019), HERO Li et al. (2020), FrozenBiLM (fine-tuned) Yang et al. (2022), and LLaMA-VQA Ko et al. (2023).
Method Transparent Full Clips TVQA
@@ 6.1 QA Evaluation
STAGE No Yes 70.5
HERO No No 74.2
FrozenBiLM No No 82.0
LLaMA-VQA No No 82.2
@@
FrozenBiLM No Yes 26.3
SeVILA No Yes 38.2
VideoChat2 No Yes 40.6
${}^{\text{\textdaggerdbl}}$ Yes Yes 44.9
TV-TREES Yes Yes 49.4
@@
Table 1: Table comparing vision-text understanding models on qualitative criteria and the TVQA benchmark. Experiment results suggest that TV-TREES and TV-TREES with text input only (${}^{\text{\textdaggerdbl}}$) outperform existing zero-shot methods on full clips. Competing method results are taken from their respective papers except for FrozenBiLM*, which we re-run on our validation subset with full clips as input. (On ground truth clip fragments, FrozenBiLM reports 59.7% accuracy).
Ablations
Existing work notes that multimodal models are often biased toward the text modality, relying on text data for reasoning even for video-centric questions. To assess TV-TREES, we report TVQA performance conditioned on input modality. We compare system output when it is only provided with dialogue transcripts and then when it is only provided with video frames.
Results
We report overall accuracy alongside qualitative comparisons between the approaches in Table 1. As shown in the table, TV-TREES outperforms existing zero-shot methods when using full clips. The influence of the individual modalities on TV-TREES is further illustrated through the ablation experiment results in Figure 6, which reports the % of questions which are correctly answered with complete trees and the % that are correctly answered with incomplete proofs given text, visual, and multimodal evidence. The results show that joint modality evidence improves both accuracy and correct tree completion in TV-TREES.
6.2 Tree Quality Evaluation
Setup   We randomly sample 600 completed entailment trees generated by TV-TREES from the TVQA validation split, split evenly between evidence modality (text vs. multimodal) and tree complexity (ranging from one to seven tree nodes). We evaluate these sampled trees using the automatic GPT-4 approach as described in Section 5.2. We then sample 200 proofs from this set (evenly distributed across modalities and complexity) and evaluate this set with human annotators from Amazon Mechanical Turk as described in Section 5.1. For human annotations, we identify careful annotators through a preliminary pilot task where each annotator’s work is scored by hand, and only high-scoring annotators are invited to annotate the full trees. More information regarding crowdsourcing is included in Appendix C.
Figure 4: Ablation experiment results comparing TV-TREES performance on TVQA using only dialogue evidence, only visual evidence, and both modalities. We report the % of questions answered correctly with completed trees and the % answered correctly overall.
Figure 5: An example of a correct completed entailment tree produced with TV-TREES using text evidence (shown below the leaf nodes). Using our tree quality evaluation system detailed in Section 6.2, the tree earns perfect scores for acceptability (the dialogue entails the sub-hypotheses), relevance and distinctness (all child hypotheses help prove the parent and are distinct), and sufficiency (there is no information lost).
Results
We report results in Table 2. For comparison, we include a high-quality tree produced by TV-TREES in Figure 5. Generally, there is a close alignment between the machine scores and human scores, but GPT-4 tends to score the text-only trees more harshly than humans, and the multimodal trees more leniently. This is shown primarily in the resulting acceptability scores, and more moderately in the sufficiency scores. GPT-4 rated relevance scores more leniently for both modalities, which may stem from differences in human interpretations of the task instructions. Distinctness scores are almost identical between the two methods.
We find that the majority of error stems from acceptability issues. According to human evaluations, the vision module produces lower-quality inferences than the textual modules do. This is not surprising, as we are able to include additional entailment filters for the textual reasoning steps to remove lower-quality predictions before constructing the final entailment trees, whereas we do not have similar methods in place for visual inference.
6.3 Qualitative Analysis
Setup
Finally, we sample a set of 120 complete but incorrect entailment trees produced by TV-TREES on the TVQA dataset and analyze them to diagnose common error patterns. We find 8 main error classes, described below and in Table 3.
Trees Accept Relev Distinct Suffic Score
@@ 6.3 Qualitative Analysis
Table 2: Entailment tree quality evaluations using human and LLM critics, reporting mean qualia scores alongside total score. We partition results by modality: Trees using text evidence only, trees that use both modalities, and both groups combined (all). As shown, tree scores largely suffer due to acceptability, highlighting the difficulty of extracting high-level inferences from dialogue and ambiguous video.
Failure Modes
Visual reasoning errors are common among erroneous proofs, especially ones involving colors and character identification (we do not implement a character identification module in TV-TREES, so this error class is not surprising). Hallucinated text inferences is another common error class, occurring in the “evidence retrieval” module (Section 4.3). Another common issue is the system ignoring negation in NL text: As documented in existing work Hosseini et al. (2021), language models and classifiers often have difficulty recognizing negation in sentences. This can lead to specific entailment misclassifications. Finally, we notice that in some cases, the dataset question wording is difficult to interpret either due to coreference ambiguity or grammatical issues.
Results
The distribution of error types among the sample set, reported in Table 3, reflects and helps to explain the tree quality evaluation results reported in Table 2: The most errors occur due to the acceptability of the produced evidence, as visual reasoning errors, hallucinated text inferences, and character and color identification account for 52% of the tree errors in the qualitative study. Entailment misclassification and ignoring textual negation accounts for 26% of the errors, explaining the lower sufficiency scores in Table 2.
Error Type % Modality
Visual reasoning errors 20% V
Hallucinated text inferences 19% T+V
Entailment misclassification 18% T+V
Ignoring negation in text 8% T+V
Character identification 7% V
Ambiguous QA pairs 7% T+V
Color identification 6% V
Other 15% _
Table 3: Distribution of error modes across a sample of 120 complete but incorrect entailment trees generated by TV-TREES, analyzed by hand. Some of the error modes are particularly prevalent subclasses of other error modes, for example, “color identification" could fall under “general visual reasoning errors". (V = Occurs in vision/multimodal proofs only, T = occurs in both multimodal and text-only proofs).
Figure 6: Examples of possible inference generation, entailment classification, and VQA filtering errors that illustrate the different failure mode categories identified in the qualitative analysis detailed in Section 6.3 and Table 3. Blue cells indicate vision-specific error types, and orange cells may occur in both text-only and multimodal proofs. Notably, "unclear question or answer" is not a failure of the system itself, but an artifact of the dataset used.